OpenStudy (anonymous):

Let f(x) = 4+5x and g(x)= 2x -1 Find g(g(x))

OpenStudy (across):

\[g(g(x)),\]\[g(2x-1),\]\[2(2x-1)-1,\]\[... ?\]

OpenStudy (across):

@Yahoo!, that's not right. :P What you wrote is equivalent to \(g(x)g(x)\), not \(g(g(x))\).

Parth (parthkohli):

Careful of the first reply.

OpenStudy (jesuschrist):

{5x+4=f(x), 2x=g(x)+1} There is your answer.

OpenStudy (anonymous):

across!

Parth (parthkohli):

Sat: You are late. She's been on this site for 1 week now.

OpenStudy (anonymous):

born late